OpenStudy (anonymous):

solve for x... 6/x=x/12 (, six is over x then x is over twelve)

OpenStudy (anonymous):

Assuming x is not zero. Multiply both sides by x, then you get 6 = 12x^2 => x^2 = 1/2 => x = +/- 1/sqrt(2).

OpenStudy (anonymous):

\[\frac{6}{x}=\frac{x}{12}\] \[x^2=72\] \[\sqrt{x^2}=\sqrt{72}\] \[x=\pm \sqrt{72}=\pm \sqrt{36*2}=\pm 6\sqrt{2}\]